OpenStudy (anonymous):

how do in go about integrating : 8sin^2(t+pi/12) dt

OpenStudy (amistre64):

does the pi/12 actually affect anything since its a constant?

OpenStudy (amistre64):

no ... of course not

OpenStudy (amistre64):

rewrite this into a lower power thru trig identities is my hunch

OpenStudy (anonymous):

isnt 8 a constant

OpenStudy (anonymous):

cant i take it out

OpenStudy (amistre64):

8 is yes

OpenStudy (amistre64):

the derivative that pops out of the sin^2 = 1 so thats just there to confuse you

OpenStudy (anonymous):

man, i hate the fact that i cant reply quickly on this site, it tells me to wait

OpenStudy (amistre64):

rewrite this as: u = t + C; du = dt \[\int 8sin^2(u)du\]

OpenStudy (amistre64):

that cleans up the clutter; now do you recall that sin^2(u) = \(\cfrac{1-cos(2u)}{2}\)??

OpenStudy (anonymous):

yes i know that half angle formula

OpenStudy (amistre64):

use it in place and itll make life easier then

OpenStudy (anonymous):

why is pi/12 a constatn

OpenStudy (amistre64):

becasue 3.14/12 is just another number that dissapears in the derivative so its pointless to write out while taking the derivetive :)

OpenStudy (anonymous):

so you replaced it with C in the u sub

OpenStudy (amistre64):

yes, just to keep from having to type it all in there and waste my fingers ....

OpenStudy (amistre64):

its still there in the end; just remember it is all :)

OpenStudy (anonymous):

now when i replace it at the end, do i flip those fractions

OpenStudy (amistre64):

u = t + 2 ; du = dt u = t + 13/54 ; du = dt u = t + pi/12 ; du = dt u = t + C ; du = dt

OpenStudy (anonymous):

oh i got

OpenStudy (amistre64):

\[\int 8sin^2(u)du\] \[8\int sin^2(u)du\] \[8\int (\frac{1}{2}-\frac{cos(2u)}{2})du\] \[8(\int \frac{1}{2}du-\int \frac{cos(2u)}{2})du\] \[8\left(\ \frac{1}{2}\int du-\frac{1}{2}\int cos(2u)\right)du\]

OpenStudy (anonymous):

oh man i got lost

OpenStudy (amistre64):

lol .... most of that is done in the head

OpenStudy (anonymous):

i know but , in the end during computation, i was usppose to get a +9 at the end

OpenStudy (amistre64):

\[\int8\frac{1-cos(2u)}{2}du\] \[\int\frac{8}{2}(1-cos(2u))du\] \[4\int(1-cos(2u))du\]

OpenStudy (amistre64):

a +9?

OpenStudy (anonymous):

yeah, i dont know either. this is an initial value problem

OpenStudy (anonymous):

where s(0)=8

OpenStudy (amistre64):

\[4\int(1-cos(2u))du\] \[4u -2sin(2u)+C\] \[4u -2sin(2(t+\frac{pi}{12}))+C\] \[4u -2sin(2t+\frac{pi}{6})+C\] oh... initial value then, ok

OpenStudy (amistre64):

i forgot a "u" :) \[4t +\frac{pi}{3} -2sin(2t+\frac{pi}{6})+C\]

OpenStudy (anonymous):

how come you didnt replace the u in 4u

OpenStudy (amistre64):

now when t = 0 we get: pi/3 -2 sin(30) + C = 8

OpenStudy (amistre64):

i forgot, im old and senile :)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but how do we get 9?

OpenStudy (amistre64):

pi/3 -2/2 + C = 8 pi/3 + C = 9 C = 9 - pi/3 if i did it right

OpenStudy (anonymous):

hmm, let me go over it ans see what i get, thanks for the help

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+8sin^2%28t%2Bpi%2F12%29+dt says i got it good

OpenStudy (anonymous):

no, i believe you did it right, i just got to get a hold of it myself :)

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

hey bro, i got it